OpenStudy (anonymous):

A 0.280kg ball makes an elastic head on collision with a second ball initially at rest.The second ball moves off with half the original speed of the first ball. what is the mass of a second ball? what fraction of the original kinetic energy gets transfered to the second ball.?

OpenStudy (mos1635):

m1*V=M1*V1+M2*V2 (conservation of momentum) M1*(V-V1)=M2*V2 (1) 1/2*M1*V^2=1/2*M1*V1^2 + 1/2*M2*V2^2 M1(V^2-V1^2)=M2*V2^2 M1*(V-V1)*(V+V1)=M2*V2^2 divide with (1) V+V1=V2 but V2=V/2 so V1=-V/2 aplay that to (1) M1*3V/2=M2*V/2 m2=3m1

OpenStudy (anonymous):

thank you